Assume Cole Corporation originally issued 300 shares of $50 par convertible preferred stock at $110 per share on December 31. Required: If each preferred share may be converted into 6 shares of S8 par common stock and all the shares are converted, what journal entry wil Cole make to record the conversion?PAGE 1 GENERAL JOURNAL DATE ACCOUNT TITLE POST. REF DEBIT CREDIT Dec. 31 Preferred Stock Common Stock Additional Paid-in Capital from Preferred Stock Conversion Additional Paid-in Capital on Prefered Stock 15,000.0041 Inventory 12 Presoid Incuanax 181 Eq-ipmen 000 Cost af Goods 8o d Insuran Enparsa 511 21 Sanes Expens 532 Esd Delt Expense 340 Irtcret Epene 231 Ealais Payable 550 Male Epor 251 Income Ts Payabe 10 Income Tc Espens 11 Comman Stack 320 Adational Faid-in Capital on Comman Etock 331 Ratained Eaming 15.000 00 Cannmon Stock din fro m Preferred Stock Conmesian

I need help finding the remaining amounts for debit and credit. Everything I put is wrong. Please show how to get answer so I can understand! Thank you!
Answer #1

Ans:

The Journal Entry after Conversion:

Date Accounting Entry and Explanation Debit Credit
Preferred Stock(Par Value of $50) $15000
Additional Paid In capital $18000
          Common Stock $14400
           Additional Paid In Capital from Stock Conversion $18600

Note:

Preferred Stock 300 shares *$50=$15000

Additional Paid In capital=300*60=$18000

Common Stock Conversion=300 shares * 6=1800*8=14400

Additional Paid In capital from stock conversion=15000+18000-14400=18600